An electron starts from rest 44.0 cm from a fixed point charge with -0.108 nC .

How fast will the electron be moving when it is very far away?
Answer #1

The electric potential at a distance r is

V=KQ/r =(9*10^9)(0.108*10^-9)/0.44

V=2.209 V

since

KE =-qV

(1/2)mv^2 =-qV

(1/2)*(9.11*10^-31)*v^2 =-(-1.6*10^-19)*2.209

v=8.81*10^5 m/s